# Non-variational Quantum Combinatorial Optimisation

### v1

@@ Abstract
This paper introduces a non-variational quantum algorithm designed to solve a wide range of combinatorial optimisation problems. The algorithm leverages an engineered interference process achieved through repeated application of two unitaries; one inducing phase-shifts dependent on objective function values, and the other mixing phase-shifted probability amplitudes via a continuous-time quantum walk (CTQW) on a problem-specific graph. The algorithm’s versatility is demonstrated through its application to various problems, namely those for which solutions are characterised by either a vector of binary variables, a vector of non-binary integer variables, or permutations (a vector of integer variables without repetition). An efficient quantum circuit implementation of the CTQW for each of these problem types is also discussed. A penalty function approach for constrained problems is also introduced, including a method for optimising the penalty function. The algorithm’s performance is demonstrated through numerical simulation for randomly generated instances of the following problems (and problem sizes): weighted maxcut (18 vertices), maximum independent set (18 vertices), k-means clustering (12 datapoints, 3 clusters), capacitated facility location (12 customers, 3 facility locations), and the quadratic assignment problem (9 locations). For each problem instance, the algorithm finds a globally optimal solution with a small number of iterations.
@@ I Introduction
On the other hand, performing the CTQW over a graph that incorporates the underlying problem structure enables exploitation of this structure, leading to speedups significantly larger than that of Grover’s algorithm. This paper introduces an algorithm that does exactly that, the non-variational quantum walk-based optimisation algorithm (non-variational QWOA). By closely analysing the CTQW and studying the statistics of objective function values as distributed over the mixing graph, it is possible to understand and design the algorithm’s interference process. This understanding helps with the intelligent selection of parameters, namely those that control the walk times and magnitudes of applied phase-shifts, thus removing the need for a variational procedure. In addition, a clear understanding of the interference process informs the design of the mixing graph, allowing for generalisation to problems with diverse and non-binary structures. Effective penalty functions can also be designed to embed problem constraints where necessary, further improving versatility.
This paper formally introduces the non-variational QWOA and focuses on its application to various combinatorial optimisation problems. Efficient circuit implementations of CTQWs for binary, non-binary and permutation problems are discussed, after which the algorithm’s performance is analysed for various combinatorial optimisation problems via numerical simulation results. A more detailed analysis of the algorithm, the general interference process, and the problem specific CTQWs is included in a separate paper [21].
@@ II-A Definitions
Consider a combinatorial optimisation problem with a solution space $S$ containing $N$ feasible solutions ${\bm{x}=(x_{1},x_{2},...,x_{n})}$ each composed of decision variables, ${x_{j}\in\{0,1,...,k-1\}}$. The solution space may contain all possible solutions of this form, of which there are $N=k^{n}$. Alternatively, it may be restricted to solutions which satisfy some constraint, such as in the case of permutation-based problems where $k=n$ and $N=n!$. The solutions to the problem can be encoded in the computational basis states of a quantum computer by allocating to each variable a sub-register containing $\lceil\log_{2}k\rceil$ qubits, such that $\bm{x}$ is represented by the solution state,
@@
where |$\ket{x_{j}}$ is the computational basis state of the $j^{\text{th}}$ sub-register which directly encodes (as a binary number) the decision variable $x_{j}$.
@@
The non-variational QWOA is designed to find optimal or near-optimal solutions to the combinatorial optimisation problem, with respect to maximisation or minimisation of an objective function $f(\bm{x})$. The interference process responsible for amplifying the measurement probability of optimal and near-optimal solution states is driven by repeated applications of two unitary operations. The first of these is the phase-separation unitary,
@@
where $Q$ is a diagonal operator such that ${Q\ket{\bm{x}}=f(\bm{x})\ket{\bm{x}}}$ and hence we define this more functionally as the phase-separation unitary, as it applies a phase shift to each solution state, proportional to the associated objective function value,
@@
where $A$ is the adjacency matrix which defines the mixing graph’s structure. This unitary is referred to as the mixing unitary, or mixer for short, since it acts to distribute probability amplitudes between the solution states. The efficient implementation of this mixer is discussed in Section III.
@@ II-B The Mixing Graph
It may often be the case (as in the mixer corresponding with Fig. 1) where the mixing unitary is applied over a quantum register with a Hilbert space larger than the feasible solution space. In this case, the adjacency matrix defines a disconnected graph with several components (isolated subgraphs). The relevant subgraph connecting just the feasible solution states is one of the components, where the others can be ignored. This is because the continuous-time quantum walk distributes probability amplitude only between connected vertices/states, such that the feasible solution states form an invariant subspace of the mixer.
@@ II-C Necessary Conditions
where $D$ is the diameter of the graph, $\phi(t)$ and $r_{\bm{x}}(t)$ are real-valued functions which are positive for sufficiently small $t$ and $\phi(t)<\pi$. In other words, probability amplitude from an initial vertex is distributed to other vertices such that the complex phase of the distributed probability amplitudes is proportional to their distance from the initial vertex.
@@
$\left(\mu_{h\bm{x}}-f(\bm{x})\right)\approx-m_{h}\left(f(\bm{x})-\mu\right),$ (8)
where $\mu_{h\bm{x}}$ is defined as the mean objective function value of solutions contained in $h_{\bm{x}}$, $\mu$ is the mean objective function value of all solutions in $S$, and $m_{h}$ is a positive constant. Secondly, the constant of proportionality $m_{h}$ should increase monotonically with increasing $h$ (up to a distance which is a considerable fraction of the graph’s diameter $D$). Examples of this relationship are plotted in section IV (Fig. 3691215).
@@ II-D The Amplified State
|$\ket{\gamma,t}=\left[\prod_{r=0}^{p-1}U_{M}\left(t_{r}\right)U_{Q}\left(\frac{\pm\gamma_{r}}{\sigma}\right)\right]\ket{s},$ (9)
where $\gamma$ and $t$ are classically controlled positive-valued parameters that determine the applied phase separations and the mixing times for each iteration. Specifically, the values of $\gamma_{r}$ increase linearly over the domain $[\beta\gamma,\gamma]$, while the mixing times $t_{r}$ decrease linearly over the domain $[\beta t,t]$. These increasing and decreasing profiles seem to indicate a connection to the adiabatic algorithm, though they are independently motivated, as discussed in more detail in [21]. Here, $\pm$ accounts for whether the goal is maximisation ($+$) or minimisation (--), and $\sigma$ is the standard deviation of objective function values, controlling for variation across problems and problem instances such that appropriate values of $\gamma$ are consistently of order $\gamma\approx 1$. An approximate value for $\sigma$ is sufficient (acquired through random sampling, for instance). An appropriate value of $\beta$, where $10<\beta<$ 1, accounts for the nature of the problem, or more specifically, the distribution of $m_{h}$ in (8). A value of $\beta=0.4$ is appropriate for many of the problems presented in this paper, while for others, a value closer to zero ($\beta=\frac{1}{p}$) is appropriate. For a particular problem type, this would only need to be determined once.
@@ II-E Repeated State Preparation and Measurement
Subject to the conditions described above, and allowing a sufficient number of iterations (polynomial in problem size), appropriate values of $\gamma$ and $t$ significantly increase the measurement probability of globally optimal and near-optimal solutions within the amplified state |$\ket{\gamma,t}$ such that a globally optimal or near-optimal solution is exceedingly likely to be measured following repeated state preparation and measurement of |$\ket{\gamma,t}$. Optionally, during the repeated state preparation and measurement, the values of $\gamma$ and $t$ can be tuned in order to improve the expectation value of the objective function.
Without prior experience applying the algorithm to a particular problem/problem size, effective values of $\gamma$ and $t$ may not be known. In this case, the locally optimal values of $\gamma$ and $t$ can be determined via gradient descent/ascent initialised at $\gamma\approx 1$, and a sufficiently small value of $t$ (dependent on the mixer). This 2-dimensional optimisation seeks only a local extrema (that closest to the origin) and requires a number of state preparations which is constant with regards to problem size $n$ and iteration count $p$. Hence, we suggest this approach, when it is necessary, is best characterised as non-variational.
The best solution measured during the process of repeated state preparation and measurement is taken as the solution, either exact or approximate, to the optimisation problem.
@@ II-F Constrained Problems
There are two ways that problem constraints can be dealt with in this framework. The first approach is to restrict the space of feasible solutions $S$ to include just those solutions which are valid and do not violate the constraints. This approach is only viable if the space of valid solutions can be characterised independently of the problem instance, i.e. just in relation to solution structure. One such example is the permutation constraint, i.e. the requirement that there be no repetition in the values of the decision variables $x_{j}$ in $\bm{x}$. Of course, this approach is only effective if an efficient mixer can be implemented, one that connects only feasible solutions and satisfies the necessary condition in (7).
The other approach is more akin to that for soft constraints in classical optimisation, where the space of feasible solutions includes valid and invalid solutions, and the constraints are enforced via additional terms in the objective function, i.e. by penalising invalid solutions. This approach is often necessary, as it is frequently the case that the space of valid solutions cannot be predetermined. That is, the validity of any specific solution is problem instance specific, and further, validity can only be determined by inspection of individual solutions, which makes it impossible to design an appropriate mixing graph connecting only valid solutions.
With regards to the penalty/soft-constraint method, we suggest that for each constraint, the penalty should include a variable term in the case where it makes sense to quantify the extent to which a constraint has been violated. We also suggest that a constant term may be appropriate, to ensure that solutions which violate a constraint by only a small amount are sufficiently penalised. A third term which corrects for bi-modality introduced by the penalty function may also improve performance, as discussed in more detail in [21]. Further, we suggest that the penalty function should be parameterised such that each of the terms is controlled by a coefficient, i.e. the objective function is expressed as $f(\bm{x})_{\bm{\lambda}}$, where ${\bm{\lambda}=(\lambda_{1},\lambda_{2},...)}$ are the coefficients for each of the terms in the penalty function.
Similar to the $\gamma$ and $t$ parameters, appropriate penalty parameters $\bm{\lambda}$ may be known from prior experience. If not, and in any case, the penalty parameters can be tuned during the repeated state preparation and measurement process. The general approach would be as follows;
@@
    Select fixed penalty parameters $\bm{\lambda}_{F}$ which adequately penalise the invalid solutions, such that the globally optimal values of $f(\bm{x})_{\bm{\lambda}_{F}}$ belong to valid solutions.
@@
    Define a second objective function, parameterised by the tunable set $\bm{\lambda}_{T}$. This objective function is used for the phase-separation unitary.
@@
    Through repeated state preparation and measurement of the amplified state, use gradient ascent/descent to tune the parameters $\{\gamma,t,\bm{\lambda}_{T}\}$ (initialised with $\bm{\lambda}_{T}=\bm{\lambda}_{F}$) so as to optimise the expectation value of $f(\bm{x})_{\bm{\lambda}_{F}}$. In other words, optimise ${}_{\bm{\lambda}_{T}}\!\bra{\gamma,t}\hat{f}_{\bm{\lambda}_{F}}\ket{\gamma,t}_{\bm{\lambda}_{T}}$, where |$\ket{\gamma,t}_{\bm{\lambda}_{T}}$ is the amplified state prepared using $f(\bm{x})_{\bm{\lambda}_{T}}$.
@@ II-G The Interference Process
The interference process is explored and analysed in detail in [21]. However, it can be understood loosely by referring to the two necessary conditions (7) and (8). When the mixing unitary distributes probability amplitude from one solution state to another, it induces a phase-shift proportional to the distance between the respective vertices on the mixing graph. Due to the nature of the CTQW, for a sufficient walk time, any particular solution state |$\ket{\bm{u}}$ receives probability amplitudes from all other solution states. This exponentially large number of individually contributing probability amplitudes is capable of producing significant constructive interference. However, under the action of the mixer alone, the probability amplitudes largely destructively interfere, due to distance dependent phase rotations applied to the contributions from each subset $h_{\bm{u}}$.
However, the phase-separation unitary rotates the phases of all contributions dependent on their associated objective function values. Due to the monotonically varying mean objective function values in each subset $h_{\bm{u}}$, the phase-separation unitary rotates the resultant probability amplitude from each subset differently, dependent on distance $h$. The distance-dependent relative phases induced by the phase-separation unitary are able to offset those induced by the mixing unitary, so as to bring the contributing probability amplitudes more or less in-phase, producing constructive or destructive interference. A closer analysis in [21] reveals why the phase separations should start small and increase, and why the mixing times should decrease, but the general interference process remains consistent throughout the iterations.
@@ III A Few Important Mixers
This section addresses the mixing unitary and its efficient implementation for three cases: binary problems ($k=2$), integer problems ($k>2$) and permutation problems ($k=n$, no repetition). Note that the action of each of these mixers on an arbitrary solution state can be shown to satisfy the necessary condition in (7), as demonstrated in [21].
III-A Binary Mixer
Consider a combinatorial problem consisting of $n$ binary decision variables ${x_{j}\in\{0,1\}}$, where solutions are characterised by vectors ${\bm{x}=(x_{1},x_{2},...,x_{n})}$ and the space of feasible solutions $S$ contains all such vectors, such that $N=2^{n}$. Each binary variable is naturally encoded in the computational basis state of a single qubit, so that the solution states are the computational basis states of an $n$ qubit register. The equal superposition is efficiently prepared,

### v2

@@ Abstract
This paper introduces a non-variational quantum algorithm designed to solve a wide range of combinatorial optimisation problems, including constrained and non-binary problems. The algorithm leverages an engineered interference process achieved through repeated application of two unitaries; one inducing phase-shifts dependent on objective function values, and the other mixing phase-shifted probability amplitudes via a continuous-time quantum walk (CTQW) on a problem-specific graph. The algorithm’s versatility is demonstrated through its application to various problems, namely those for which solutions are characterised by either a vector of binary variables, a vector of non-binary integer variables, or permutations (a vector of integer variables without repetition). An efficient quantum circuit implementation of the CTQW for each of these problem types is also discussed. A penalty function approach for constrained problems is also introduced, including a method for optimising the penalty function. The algorithm’s performance is demonstrated through numerical simulation for randomly generated instances of the following problems (and problem sizes): weighted maxcut (18 vertices), maximum independent set (18 vertices), k-means clustering (12 datapoints, 3 clusters), capacitated facility location (12 customers, 3 facility locations), and the quadratic assignment problem (9 locations). For each problem instance, the algorithm finds a globally optimal solution with a small number of iterations.
@@ I Introduction
On the other hand, performing the CTQW over a graph that incorporates the underlying problem structure enables exploitation of this structure, leading to speedups significantly larger than that of Grover’s algorithm. This paper introduces an algorithm that does exactly that, the non-variational quantum walk-based optimisation algorithm (non-variational QWOA). By closely analysing the CTQW and studying the statistics of objective function values as distributed over the mixing graph, it is possible to understand and design the algorithm’s interference process. This understanding helps with the intelligent selection of parameters, namely those that control the walk times and magnitudes of applied phase-shifts, thus removing the need for a computationally expensive variational procedure. In addition, a clear understanding of the interference process informs the design of the mixing graph, allowing for generalisation to problems with diverse and non-binary structures. Effective penalty functions can also be designed to embed problem constraints where necessary, further improving versatility.
This paper formally introduces the non-variational QWOA and focuses on its application to various combinatorial optimisation problems. Efficient circuit implementations of CTQWs for binary, non-binary and permutation problems are discussed, after which the algorithm’s performance is analysed via numerical simulation results. Detailed benchmarking will form part of future work and a more detailed analysis of the algorithm, particularly the general interference process, and the problem specific CTQWs is included in a separate paper [21].
@@ II-A Definitions
Consider a combinatorial optimisation problem with a solution space $S$ containing $N$ feasible solutions ${\bm{x}=(x_{1},x_{2},...,x_{n})}$ each composed of $n$ decision variables, ${x_{j}\in\{0,1,...,k-1\}}$. The solution space may contain all possible solutions of this form, of which there are $N=k^{n}$. Alternatively, it may be restricted to solutions which satisfy some constraint, such as in the case of permutation-based problems where $k=n$ and $N=n!$. The solutions to the problem can be encoded in the computational basis states of a quantum computer by allocating to each variable a sub-register such that $\bm{x}$ is represented by the solution state,
@@
where |$\ket{x_{j}}$ is the computational basis state of the $j^{\text{th}}$ sub-register which directly encodes the decision variable $x_{j}$. If using a binary encoding, each sub-register will be assigned $\lceil\log_{2}k\rceil$ qubits, whereas a one-hot encoding will use $k$ qubits. The binary encoding is more space efficient and lends itself to a more efficient implementation of the mixing unitary. However, the one-hot encoding may enable a more efficient implementation of the phase-separation unitary.
@@
The non-variational QWOA is designed to find optimal or near-optimal solutions to the combinatorial optimisation problem, with respect to optimisation of an objective function $f(\bm{x})$. The interference process responsible for amplifying the measurement probability of optimal and near-optimal solution states is driven by repeated applications of two unitary operations. The first of these is the phase-separation unitary,
@@
where $Q$ is a diagonal operator such that ${Q\ket{\bm{x}}=f(\bm{x})\ket{\bm{x}}}$. This unitary applies a phase shift to each solution state, proportional to the associated objective function value,
@@
where $A$ is the adjacency matrix which defines the mixing graph’s structure. This unitary is referred to as the mixing unitary, or mixer for short, since it acts to distribute probability amplitudes between the solution states. The mixing graph is customised to suit the underlying structure of the problem to be solved. The efficient implementation of this mixer is discussed in Section III.
@@ II-B The Mixing Graph
It may often be the case where the mixing unitary is applied over a quantum register with a Hilbert space larger than the feasible solution space (as in the mixer corresponding with Fig. 1 where the Hilbert space of each sub-register has dimension larger than $k=3$). In this case, the adjacency matrix defines a disconnected graph with several components (isolated subgraphs). The relevant subgraph connecting just the feasible solution states, which we refer to as the mixing graph, is just one of the components, where the others can be ignored. This is because the continuous-time quantum walk distributes probability amplitude only between connected vertices/states, such that the feasible solution states form an invariant subspace of the mixer.
@@ II-C Necessary Conditions
where $D$ is the diameter of the graph and (for sufficiently small $t$) $r_{\bm{x}}$ and $\phi$ are positive real-valued functions, also $0<\phi(t)<\pi$. In other words, probability amplitude from an initial vertex is distributed to other vertices such that the complex phase of the distributed probability amplitudes is proportional to their distance from the initial vertex.
@@
$\left(\mu_{h\bm{x}}-f(\bm{x})\right)\approx-\alpha_{h}\left(f(\bm{x})-\mu\right),$ (8)
where $\mu_{h\bm{x}}$ is defined as the mean objective function value of solutions contained in $h_{\bm{x}}$, $\mu$ is the mean objective function value of all solutions in $S$, and $\alpha_{h}$ is a positive constant. Secondly, the constant of proportionality $\alpha_{h}$ should increase monotonically with increasing $h$ (up to a distance which is a considerable fraction of the graph’s diameter $D$). Examples of this relationship are plotted in section IV (Fig. 3691114).
@@ II-D The Amplified State
|$\ket{\gamma,t,\beta}=\left[\prod_{i=0}^{p-1}U_{M}\left(t_{i}\right)U_{Q}\left(\frac{\pm\gamma_{i}}{\sigma}\right)\right]\ket{s},$ (9)
where $\gamma$, $t$ and $\beta$ are classically controlled positive-valued parameters that determine the applied phase separations and the mixing times for each iteration. Specifically, the values of $\gamma_{i}$ and $t_{i}$ are given by explicit formulae, such that $\gamma_{i}$ increase linearly over the domain $[\beta\gamma,\gamma]$, while the mixing times $t_{i}$ decrease linearly over the domain $[\beta t,t]$, where $10<\beta<$ 1. These increasing and decreasing profiles are reminiscent of the annealing protocols for the adiabatic algorithm, though they are independently motivated. Here, $\pm$ accounts for whether the goal is maximisation ($+$) or minimisation (--), and $\sigma$ is the standard deviation of objective function values, controlling for variation across problems and problem instances such that appropriate values of $\gamma$ are consistently of order $\gamma\approx 1$. An approximate value for $\sigma$ is sufficient (acquired through random sampling, for instance).
@@ II-E Repeated State Preparation and Measurement
Subject to the conditions described above, and allowing a sufficient number of iterations (polynomial in problem size), a wide range of appropriate values for $\gamma$, $t$ and $\beta$ significantly increase the measurement probability of globally optimal and near-optimal solutions within the amplified state |$\ket{\gamma,t,\beta}$, such that a globally optimal or near-optimal solution is exceedingly likely to be measured following repeated state preparation and measurement of |$\ket{\gamma,t,\beta}$.
The total number of state preparations should be fixed, regardless of iteration count $p$ or problem size. The best solution measured during the process of repeated state preparation and measurement is taken as the solution, either exact or approximate, to the optimisation problem. Optionally, at regular intervals during the repeated state preparation and measurement, the values of $\gamma$, $t$ and $\beta$ can be updated in order to improve the amplification of optimal and near-optimal solutions, as indicated by improvement in approximated values of either the expectation value of the objective function, or a related measure, such as the Conditional Value at Risk (CVaR) [22].
The reason that the algorithm remains effective while constrained to a fixed total number of state-preparations and measurements, and the reason it is best characterised as non-variational, is because the optimal set of parameters can be determined via a fixed complexity 3-dimensional optimisation which seeks only a local extremum (that closest to the origin). In any case, even sub-optimal parameter values significantly amplify optimal and near-optimal solutions, such that the improvement/optimisation of parameters is optional.
@@ II-F Constrained Problems
There are two ways that problem constraints can be dealt with in this framework. The first approach is to restrict the space of feasible solutions $S$ to include just those solutions which are valid (do not violate the constraints). This approach is only viable if the space of valid solutions can be characterised independently of the problem instance, if it is possible to efficiently implement a mixer that connects only valid solutions and if the mixer satisfies the necessary condition in (7). An example of this approach is demonstrated in IV-E for the quadratic assignment problem.
The other approach is to include both valid and invalid solutions and for the constraints to manifest as additional terms in the objective function, which act to penalise invalid solutions. This approach is often necessary, as it is frequently the case that the space of valid solutions cannot be predetermined, in which case it is not possible to design a mixing graph which isolates and mixes between only valid solutions.
We introduce three kinds of penalty terms, the first two are designed to ensure that invalid solutions are sufficiently penalised, so that the penalised objective function is optimised by valid solutions. The first of these is a variable term which scales with the extent to which a constraint has been violated. The second is a fixed term which ensures that solutions which violate a constraint by only a small amount remain sufficiently penalised. The introduction of these first two penalty terms may adversely effect algorithm performance, due to the introduction of bimodality in the distribution of objective function values, which adversely effects conformance with condition (8). The third penalty term is introduced to correct for this bimodality and improve algorithm performance, while leaving near-optimal invalid solutions adequately penalised.
The influence of each penalty term is controlled by a positive-valued parameter, such that the penalty can be adjusted to improve performance. The penalised objective function is therefore expressed as $f(\bm{x})_{\bm{\lambda}}$, where ${\bm{\lambda}=(\lambda_{1},\lambda_{2},...)}$ are the coefficients for each of the terms in the penalty function. Similar to the $\gamma$, $t$ and $\beta$ parameters, appropriate penalty parameters $\bm{\lambda}$ may be known from prior experience. In any case, the penalty parameters can be tuned during the repeated state preparation and measurement process. The general approach is summarised as follows:
@@
    Design the penalty terms and select fixed penalty parameters $\bm{\lambda}_{F}$ which adequately penalise the invalid solutions, such that the globally optimal values of $f(\bm{x})_{\bm{\lambda}_{F}}$ belong to valid solutions.
@@
    Using the same penalty terms, define a second objective function, parameterised by tunable penalty parameters $\bm{\lambda}_{T}$. This objective function $f(\bm{x})_{\bm{\lambda}_{T}}$ is used within the phase-separation unitary.
@@
    Through repeated state preparation and measurement of the amplified state, use a classical optimistion procedure to tune the parameters $\{\gamma,t,\beta,\bm{\lambda}_{T}\}$ (initialised with $\bm{\lambda}_{T}=\bm{\lambda}_{F}$) so as to optimise the expectation value of $f(\bm{x})_{\bm{\lambda}_{F}}$. In other words, optimise ${}_{\bm{\lambda}_{T}}\!\bra{\gamma,t,\beta}\hat{f}_{\bm{\lambda}_{F}}\ket{\gamma,t,\beta}_{\bm{\lambda}_{T}}$, where |$\ket{\gamma,t,\beta}_{\bm{\lambda}_{T}}$ is the amplified state prepared using $f(\bm{x})_{\bm{\lambda}_{T}}$.
@@ II-G The Interference Process
The interference process can be understood loosely by referring to the two necessary conditions (7) and (8). When the mixing unitary distributes probability amplitude from one solution state to another, it induces a phase-shift proportional to the distance between the respective vertices on the mixing graph. Due to the nature of the CTQW, for a sufficient walk time, any particular solution state |$\ket{\bm{u}}$ receives probability amplitudes from all other solution states. This exponentially large number of individually contributing probability amplitudes is capable of producing significant constructive interference. However, under the action of the mixer alone, the probability amplitudes largely destructively interfere, due to the distance dependent phase rotations applied to the contributions from each subset $h_{\bm{u}}$.
Application of the phase-separation unitary rotates the phases of all contributions dependent on their associated objective function values. Due to the monotonically varying mean objective function values in each subset $h_{\bm{u}}$, the phase-separation unitary rotates the resultant probability amplitude from each subset differently, dependent on distance $h$. The distance-dependent relative phases induced by the phase-separation unitary are able to offset those induced by the mixing unitary, so as to bring the contributing probability amplitudes more or less in-phase, producing constructive or destructive interference. A closer analysis in [21] reveals why the phase separations should start small and increase, and why the mixing times should decrease, but the general interference process remains consistent throughout the iterations.
@@ III A Few Important Mixers
This section addresses the mixing unitary and its efficient implementation for three cases: binary problems ($k=2$), integer problems ($k>2$) and permutation problems ($k=n$, no repetition). The action of each of these mixers on an arbitrary solution state can be shown to satisfy the necessary condition in (7).
III-A Binary Mixer (Hypercube Graph)
Consider a combinatorial problem consisting of $n$ binary decision variables ${x_{j}\in\{0,1\}}$, where solutions are characterised by vectors ${\bm{x}=(x_{1},x_{2},...,x_{n})}$ and the space of feasible solutions $S$ contains all such vectors, such that $N=2^{n}$. Each binary variable is naturally encoded in the computational basis state of a single qubit (where the binary and one-hot encodings are equivalent), so that the solution states are the computational basis states of an $n$ qubit register. The equal superposition is efficiently prepared,